OpenStudy (anonymous):

Find the absolute extrema of the function on the closed interval. Use a graphing utility to verify your results. h(t)=(t-1)^2/3 [-7,2] <--interval

OpenStudy (anonymous):

For the absolute minimum I got (-7,-4). For the absolute maximum I got (2,1). The absolute maximum shown in the book is (-7,4)..did i do something wrong?

OpenStudy (rogue):

Well, your absolute min is wrong as well ;) You can figure out the min and max from thinking about how the graph looks, but since this is calculus, I guess you have to find them via differentiation :P What did you get for the first derivative of h(t)?

OpenStudy (anonymous):

see i don't know why my teacher does the derivative when i can find the Critical numbers simply through the original function. But it's 2/3(t-1)^-2/3

OpenStudy (rogue):

Yeah, that's correct. Now the critical points of the function are where the derivative equals zero or doesn't exist. Where does that occur?

OpenStudy (rogue):

the power of (t-1) in the derivative should actually be -1/3

OpenStudy (anonymous):

no it's -2/3 b/c it's minus 1 when you do product rule...and it's 1.

OpenStudy (rogue):

\[\frac {2}{3} - \frac {3}{3} = \frac {-1}{3}\] hehe ;D

OpenStudy (anonymous):

oh right -_-

OpenStudy (rogue):

Anyways, the derivative is \[h(t)= \frac {2}{3}(t−1)^{−1/3}=\frac {2}{3(t−1)^{1/3}}\] There are no instances where the derivative equals 0, but at t = 1, the derivative is undefined, so that is a critical point. Can you go on from there?

OpenStudy (anonymous):

for the most part yes, but the interval is [-7,2], what do you get when you substitute -7 in the original equation? What do you get for h(t)?

OpenStudy (rogue):

h(-7) = (1 - -7)^(2/3) = 8 ^ (2/3) = cube root of 8^2 = 4

OpenStudy (anonymous):

but it's t-1, not 1-t...so (-7-1)^(2/3)=-4 correct?

OpenStudy (rogue):

oh, yes, but either way, it's positive 4. cube root of (-8)^2 = cube root of 64 = 4 (cube root of -8)^2 = (-2)^2 = 4

OpenStudy (anonymous):

ah ok. thanks! my calculator was giving me -4 for some reason...

OpenStudy (anonymous):

and it still is...

OpenStudy (anonymous):

nvm, parenthesis issue -_-

OpenStudy (rogue):

If you plug in the critical value of t = 1, you'll get the absolute min of 0 btw.